OpenStudy (rainbow_dash):

for the expansion of (x+4)^8, find the 6th term.

OpenStudy (anonymous):

\[\dbinom{8}{5}x^34^5\]

OpenStudy (anonymous):

exponents must add to 8 \(\dbinom{8}{5}\) read "eight choose five" compute via compute \(\dbinom{8}{5}\) via \(\frac{8\times 7\times 6}{3\times 2}\) cancel first multiply last

OpenStudy (rainbow_dash):

okay so 56^3 (56 from the 8x7x6/3x2) and 4^5 multiply together for answer?

OpenStudy (rainbow_dash):

no nevermind that's wrong :/